OpenStudy (anonymous):

What is the mass of a stone moving at a speed of 15 meters/second and having a momentum of 7.5 kilogram·meters/second?

OpenStudy (mrnood):

the clue is in the units \[\frac{ \frac{ kg m }{ s } }{ \frac{ m }{ s } } =kg\] so divide the momentum by the velocity to get the mass

OpenStudy (anonymous):

do any body knows what the awnser is

OpenStudy (mrnood):

yes - I do as I said above (7 months ago) divide the momentum by the velocity to get the mass

OpenStudy (anonymous):

but whats the velocity

OpenStudy (mrnood):

well - loosely speaking - it's the speed as given in the question: "a stone moving at a speed of 15 meters/second "